OpenStudy (anonymous):

Solve the equation. 2^(1 + 2x) = 32

OpenStudy (anonymous):

\[2^{1+2x} = 32\]

OpenStudy (anonymous):

lol yup. it's a log problem, and i'm beyond horrid at it. so idk if i can help, but i'll see what i can do.

OpenStudy (anonymous):

2^(1 + 2x) = 32 raise both sides to the log of base 2 to eliminate the left side's base. (1 + 2x) = log base 2 of 32 1 + 2x = 2^5 = 32 therefore, 5 1 + 2x = 5 2x = 4 x=2

OpenStudy (anonymous):

or without logs note that \[32=2^5\] so \[1+2x=5\] and therefore \[2x=4\] \[x=2\]

OpenStudy (anonymous):

heh yup satellite did it better, thanks actually i needed that too. :x